# Compressed code: the hidden effects of quantization and distillation on programming tokens

### v1

@@ 1.1 Motivation
Programming language understanding and code generation are critical applications of LLMs. Tokenization plays a crucial role in shaping model efficiency and performance. Byte Pair Encoding (bostrom2020byte) is widely used but often introduce inefficiencies, particularly in programming languages where syntax-specific tokens must be preserved. Poor tokenization can lead to increased sequence length, higher computational cost, and reduced model interpretability.
While LLMs excel at natural language tasks, their ability to handle programming languages remains dependent on how well their tokenizers preserve syntax and semantics. Many tokenizers are optimized for general text but fail to efficiently capture code structure, leading to suboptimal code completion and generation results (see dagan2024tokenizer). Recent work has shown that tokenizer quality significantly impacts model performance across different languages and domains (chelombitko2024qtokcomprehensiveframeworkevaluating), with specialized tokenizers demonstrating substantial improvements in compression ratios and representation efficiency (see chelombitko-komissarov-2024-specialized). An emerging alternative direction explores tokenizer-free architectures: the Meta-AI Byte Latent Transformer (see pagnoni2024bytelatenttransformerpatches) demonstrates that byte-level models can match tokenization-based LLM performance at scale, suggesting that the relationship between tokenization and model capabilities may be more nuanced than previously assumed.
@@ 1.2 Background and Related Work
Despite these advances, few studies delve deeply into the granularity of token representations for programming languages. Existing research often assesses model performance via downstream tasks using traditional benchmarks like HumanEval (chen2021evaluating) or MBPP (austin2021program). However, the shift towards evaluating complex instruction following and repository-level reasoning exemplified by BigCodeBench (zhuo2025bigcodebenchbenchmarkingcodegeneration) and SWE-bench (jimenez2024swebenchlanguagemodelsresolve) - highlights the need for deeper interpretability. These modern benchmarks reveal that high pass rates do not always correlate with structural understanding, yet few studies link these performance metrics to the granularity of token representations or probability evolution.
Common model compression strategies include quantization (reducing numerical precision), distillation (teacher-student paradigms), and pruning (removing weights or neurons). While compression methods are widely studied for natural language tasks, code-focused compression remains underexplored. Preliminary evidence suggests that quantization can degrade the ability to handle rare tokens (ouyang2024lowbit), and distillation can lead to biased distributions if specialized tokens are not preserved.
@@ 2.1 Choosing coding models and tokenizers
As a starting point in our study, we selected a set of top-performing open-source language models that demonstrated the best performance in coding tasks while also being publicly available in the Chatbot Arena benchmark (chiang2024chatbot). The models included in our analysis are Qwen2.5-Coder, Qwen2.5 (hui2024qwen25), Athene-V2-Chat (nexusflow2024athenev2), DeepSeek-V2.5 (deepseek2024deepseekv2), DeepSeek-V3 (deepseek2024deepseekv3), DeepSeek-R1, and additionally DeepSeek-R1 distillations: DeepSeek-R1-Llama-70B, DeepSeek-R1-Llama-8B, DeepSeek-R1-Qwen-32B, DeepSeek-R1-Qwen-14B, DeepSeek-R1-Qwen-7B, DeepSeek-R1-Qwen-1.5B (deepseek2025deepseekr1), and Llama 3.1 (grattafiori2024llama3).
@@ 2.2 Keywords as language markers in tokenizers
Next, we considered the most popular programming languages on GitHub (githuttwo), as this metric better reflects real-world usage and developer activity. Based on recent rankings, we selected the top ten programming languages: Python, Java, Go, JavaScript, C++, TypeScript, PHP, Ruby, C, and C#. Additionally, we included Rust and React, the last is a widely used JavaScript library for building user interfaces, as it has gained significant traction in code generation tasks using large language models. To assess how well these languages and frameworks are represented in the tokenizers, we adopted the following approach. For each programming language, we compiled a set of its reserved keywords, while for React, we collected key framework-specific keywords. These sets were then used as a basis for evaluating their presence in the tokenizers.
@@ 2.7.1 Quantization Effects
For a deeper understanding of how model compression affects coding abilities, we analyzed various quantization levels of the Qwen2.5-coder-7B model. It is interesting to note that the quantized Q8_0 model is very close to the original model. Our analysis revealed several critical patterns that challenge initial assumptions about model performance. A key finding emerged in the distribution between programming keywords and special tokens. While Q2_K showed a concerning ratio with high special token probability (0.4248) compared to keyword probability (0.0902), less aggressive quantization actually improved this balance. The Q4_K_S variant, for instance, demonstrated a more favorable distribution with reduced special token probability (0.2804) and increased keyword probability (0.1231). This observation aligns with recent findings on quantization robustness (fang2025smallerweakerbenchmarking), which demonstrate that quantized models can exhibit superior resilience to noise perturbations compared to their full-precision counterparts. This shift is particularly significant because high special token probabilities indicate a model’s tendency to generate noise – repeated spaces, brackets, hashes, and other non-semantic elements that can lead to degenerate output patterns. For example, we observed tokens representing multiple repeated spaces, which could cause the model to get stuck in loops of meaningless character generation. The consistency in top programming keywords (’import’, ’package’, ’def’, ’from’) across quantization levels suggests that essential coding semantics are preserved. However, the reduced probability of special tokens in more aggressive quantization schemes might actually represent an improvement in generation quality, as it indicates a lower likelihood of producing syntactic noise.
@@ 2.9 Difference with distilled models
The comparison between original Qwen models and their distilled DeepSeek-R1 counterparts reveals striking differences in token probability distributions, suggesting significant transformations in model behavior through the distillation process. Most notably, the distilled models show lower programming keyword probabilities compared to their original counterparts. For instance, while Qwen2.5-1.5B maintains a healthy keyword probability of 0.143, its distilled version (DeepSeek-R1-Distill-Qwen-1.5B) drops to merely 0.0042 – a reduction of nearly 97%. This pattern persists across all model sizes, with the 32B variant showing a similar decrease from 0.142 to 0.0019. Even more concerning is the shift in special token probabilities. The distilled 1.5B model shows an increase in special token probability (0.7246) compared to its original version (0.125), suggesting a strong bias toward generating syntactic noise. This indicates that the distillation process might be inadvertently amplifying the model’s tendency to produce non-semantic tokens while diminishing its ability to generate meaningful programming constructs. The transformation is also evident in the models’ preferred tokens. While original Qwen-family models consistently favor semantic programming keywords (’import’, ’package’, ’from’), the distilled versions show a shift toward more basic constructs (’void’, ’int’, ’is’) and exhibit higher probabilities for formatting-related special tokens. This suggests that the distillation process, while potentially optimizing for model size and speed, might be compromising the models’ ability to generate coherent code. This degradation mirrors the theoretical findings (dohmatob2024modelcollapsedemystifiedcase) regarding ’model collapse’. Their analysis demonstrates that models trained on synthetic outputs (as in the distillation process) tend to lose variance in their probability distributions, effectively pruning the vocabulary tails and converging onto a narrow set of high-frequency tokens. It raises important considerations about the trade-offs involved in model distillation, particularly for code generation tasks where maintaining a balance between semantic and syntactic elements is crucial.

### v2

@@ 1.1 Motivation
Programming language understanding and code generation are critical applications of LLMs. Tokenization plays a crucial role in shaping model efficiency and performance. Byte Pair Encoding (Bostrom and Durrett, 2020) is widely used but often introduce inefficiencies, particularly in programming languages where syntax-specific tokens must be preserved. Poor tokenization can lead to increased sequence length, higher computational cost, and reduced model interpretability.
While LLMs excel at natural language tasks, their ability to handle programming languages remains dependent on how well their tokenizers preserve syntax and semantics. Many tokenizers are optimized for general text but fail to efficiently capture code structure, leading to suboptimal code completion and generation results (see Dagan and others 2024). Recent work has shown that tokenizer quality significantly impacts model performance across different languages and domains (Chelombitko et al. 2024), with specialized tokenizers demonstrating substantial improvements in compression ratios and representation efficiency (see Chelombitko and Komissarov 2024). An emerging alternative direction explores tokenizer-free architectures: the Meta-AI Byte Latent Transformer (see Pagnoni et al. 2024) demonstrates that byte-level models can match tokenization-based LLM performance at scale, suggesting that the relationship between tokenization and model capabilities may be more nuanced than previously assumed.
@@ 1.2 Background and Related Work
Despite these advances, few studies delve deeply into the granularity of token representations for programming languages. Existing research often assesses model performance via downstream tasks using traditional benchmarks like HumanEval (Chen and others, 2021) or MBPP (Austin and others, 2021). However, the shift towards evaluating complex instruction following and repository-level reasoning exemplified by BigCodeBench (Zhuo et al., 2025) and SWE-bench (Jimenez et al., 2024) - highlights the need for deeper interpretability. These modern benchmarks reveal that high pass rates do not always correlate with structural understanding, yet few studies link these performance metrics to the granularity of token representations or probability evolution.
Common model compression strategies include quantization (reducing numerical precision), distillation (teacher-student paradigms), and pruning (removing weights or neurons). While compression methods are widely studied for natural language tasks, code-focused compression remains underexplored. Preliminary evidence suggests that quantization can degrade the ability to handle rare tokens (Ouyang and others, 2024), and distillation can lead to biased distributions if specialized tokens are not preserved.
@@ 2.1 Choosing coding models and tokenizers
As a starting point in our study, we selected a set of top-performing open-source language models that demonstrated the best performance in coding tasks while also being publicly available in the Chatbot Arena benchmark (Chiang and others, 2024). The models included in our analysis are Qwen2.5-Coder, Qwen2.5 (Hui and others, 2024), Athene-V2-Chat (Nexusflow, 2024), DeepSeek-V2.5 (DeepSeek-AI et al., 2024a), DeepSeek-V3 (DeepSeek-AI et al., 2024b), DeepSeek-R1, and additionally DeepSeek-R1 distillations: DeepSeek-R1-Llama-70B, DeepSeek-R1-Llama-8B, DeepSeek-R1-Qwen-32B, DeepSeek-R1-Qwen-14B, DeepSeek-R1-Qwen-7B, DeepSeek-R1-Qwen-1.5B (DeepSeek-AI et al., 2025), and Llama 3.1 (Grattafiori and others, 2024).
@@ 2.2 Keywords as language markers in tokenizers
Next, we considered the most popular programming languages on GitHub (Beuke, 2023), as this metric better reflects real-world usage and developer activity. Based on recent rankings, we selected the top ten programming languages: Python, Java, Go, JavaScript, C++, TypeScript, PHP, Ruby, C, and C#. Additionally, we included Rust and React, the last is a widely used JavaScript library for building user interfaces, as it has gained significant traction in code generation tasks using large language models. To assess how well these languages and frameworks are represented in the tokenizers, we adopted the following approach. For each programming language, we compiled a set of its reserved keywords, while for React, we collected key framework-specific keywords. These sets were then used as a basis for evaluating their presence in the tokenizers.
@@ 2.7.1 Quantization Effects
For a deeper understanding of how model compression affects coding abilities, we analyzed various quantization levels of the Qwen2.5-coder-7B model. It is interesting to note that the quantized Q8_0 model is very close to the original model. Our analysis revealed several critical patterns that challenge initial assumptions about model performance. A key finding emerged in the distribution between programming keywords and special tokens. While Q2_K showed a concerning ratio with high special token probability (0.4248) compared to keyword probability (0.0902), less aggressive quantization actually improved this balance. The Q4_K_S variant, for instance, demonstrated a more favorable distribution with reduced special token probability (0.2804) and increased keyword probability (0.1231). This observation aligns with recent findings on quantization robustness (Fang et al., 2025), which demonstrate that quantized models can exhibit superior resilience to noise perturbations compared to their full-precision counterparts. This shift is particularly significant because high special token probabilities indicate a model’s tendency to generate noise – repeated spaces, brackets, hashes, and other non-semantic elements that can lead to degenerate output patterns. For example, we observed tokens representing multiple repeated spaces, which could cause the model to get stuck in loops of meaningless character generation. The consistency in top programming keywords (’import’, ’package’, ’def’, ’from’) across quantization levels suggests that essential coding semantics are preserved. However, the reduced probability of special tokens in more aggressive quantization schemes might actually represent an improvement in generation quality, as it indicates a lower likelihood of producing syntactic noise.
@@ 2.9 Difference with distilled models
The comparison between original Qwen models and their distilled DeepSeek-R1 counterparts reveals striking differences in token probability distributions, suggesting significant transformations in model behavior through the distillation process. Most notably, the distilled models show lower programming keyword probabilities compared to their original counterparts. For instance, while Qwen2.5-1.5B maintains a healthy keyword probability of 0.143, its distilled version (DeepSeek-R1-Distill-Qwen-1.5B) drops to merely 0.0042 – a reduction of nearly 97%. This pattern persists across all model sizes, with the 32B variant showing a similar decrease from 0.142 to 0.0019. Even more concerning is the shift in special token probabilities. The distilled 1.5B model shows an increase in special token probability (0.7246) compared to its original version (0.125), suggesting a strong bias toward generating syntactic noise. This indicates that the distillation process might be inadvertently amplifying the model’s tendency to produce non-semantic tokens while diminishing its ability to generate meaningful programming constructs. The transformation is also evident in the models’ preferred tokens. While original Qwen-family models consistently favor semantic programming keywords (’import’, ’package’, ’from’), the distilled versions show a shift toward more basic constructs (’void’, ’int’, ’is’) and exhibit higher probabilities for formatting-related special tokens. This suggests that the distillation process, while potentially optimizing for model size and speed, might be compromising the models’ ability to generate coherent code. This degradation mirrors the theoretical findings (Dohmatob et al., 2024) regarding ’model collapse’. Their analysis demonstrates that models trained on synthetic outputs (as in the distillation process) tend to lose variance in their probability distributions, effectively pruning the vocabulary tails and converging onto a narrow set of high-frequency tokens. It raises important considerations about the trade-offs involved in model distillation, particularly for code generation tasks where maintaining a balance between semantic and syntactic elements is crucial.